OpenStudy (anonymous):

A bag contains 3 red marbles, 4 blue marbles, and 5 white marbles. What is the probability of pulling out a white marble, and then a blue marble if none are replaced? A.5/33 B.9/23 C.5/36 D.9/121

OpenStudy (anonymous):

\[=\frac{5}{12}*\frac{4}{11}=\frac{9}{132}=\frac{3}{44}\]

OpenStudy (anonymous):

ooops...20/132=5/33 lol

OpenStudy (anonymous):

apparently I think 5*4=9 hahaha